The random variable X is exponentially distributed with mean 3. Find P(X > t + 3|X > t)
where t is any positive real number.
1
Expert's answer
2021-08-16T17:13:38-0400

Gicen "X\sim Exp(\lambda), \mu=3"


"\lambda=\dfrac{1}{\mu}=\dfrac{1}{3}"

By the Memoryless Property:


"P(X>t+r|X>t)=P(X>r), r\geq0, t\geq0"

Then


"P(X>t+3|X>t)=P(X>3)"

"=e^{-3({1 \over 3})}=e^{-1}\approx0.3679"